OpenStudy (anonymous):

Using complete sentences, explain how to completely factor 2x^2 - 8x - 42.

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

Lol jk <3

OpenStudy (anonymous):

divide by 2 first

OpenStudy (anonymous):

Find the GCF which would be 2. Put it on the outside of the problem.

OpenStudy (anonymous):

Find the x intercepts f(x)=2x^2 +8x-42 --- Let y = 0 2x^2 + 8x - 42 = 0 2(x^2 + 4x - 21) = 0 Factor: 2(x+7)(x-3) = 0 x = -7 or x = 3

OpenStudy (anonymous):

x=7, x=-3

OpenStudy (anonymous):